Witter House is a calendar-year firm with 410 million common shares outstanding throughout 2018 and 2019. As part of its executive compensation plan, at January 1, 2017, the company had issued 50 million executive stock options permitting executives to buy 50 million shares of stock for $15 within the next eight years, but not prior to January 1, 2020. The fair value of the options was estimated on the grant date to be $3 per option.

In 2018, Witter House began granting employees stock awards rather than stock options as part of its equity compensation plans and granted 25 million restricted common shares to senior executives at January 1, 2018. The shares vest four years later. The fair value of the stock was $20 per share on the grant date. The average price of the common shares was $20 and $25 during 2018 and 2019, respectively.

The stock options qualify for tax purposes as an incentive plan. The restricted stock does not. The company's net income was $260 million and $270 million in 2018 and 2019, respectively. Its income tax rate is 40%.

Required:

1. Compute basic and diluted earnings per share for Witter House in 2018.

2. Compute basic and diluted earnings per share for Witter House in 2019.

(For all requirements, do not round intermediate calculations. Enter your answers in millions rounded to 2 decimal places (i.e., 10,000,000 should be entered as 10.00).)
Answer #1

Basis EPS = Net income after tax / Total no. of shares outstanding

Net income after tax = Net income - tax on income

2018 = $260 - ($260 * 40%) = $156

2019 = $270 - ($270 * 40%) = $162

Basis EPS = Net income after tax / Total no. of shares outstanding

2018 = $156 / 410 = 0.38 EPS

2019 = $162 / 410 = 0.39 EPS

1. Net ncome after tax for 2018 = $260 - ($260 * 40%) = $156

Diluted shares = 410 + 25 = 435

Diluted EPS = $156/ 435 = 0.36 EPS

2.  Net ncome after tax for 2019 = $270 - ($270 * 40%) = $162

Diluted shares = 410 + 25 = 435

Diluted EPS = $162 / 435 = 0.37 EPS